# A Curious Case of Remarkable Resilience to Gradient Attacks via Fully Convolutional and Differentiable Front End with a Skip Connection

### v1

@@ 3.3 Standard and Adversarial Training
To train models resistant to adversarial attacks Szegedy et al. (2014) and Goodfellow et al. (2015) proposed to expand or replace the original test with adversarial examples generated by the model. This general idea has several concrete realizations, which are broadly termed as adversarial training. As noted by Goodfellow et al. (2015), “the adversarial training procedure can be seen as minimizing the worst case error when the data is perturbed by an adversary.” Using a popular formalization of this statement due to Madry et al. (2018), an adversarial training objective can be written as the following minimax optimization problem:
@@
The inner maximization problem in Equation 2 corresponds to generation of adversarial example for a given data point and a model. This problem is generally intractable: In practice it is solved approximately using methods such as projected gradient descent (Biggio et al., 2013; Madry et al., 2018). In Section 4.2 we describe several approaches for finding adversarial examples (including PGD), which are used in our experiments.
Figure 1: Denoising front end (DnCNN) architecture Zhang et al. (2017)): Note the shortcut connection.
4 Methods
@@ 4.1 Model Architectures and Training
  • Standard, i.e., non-robust models, ,i.e., which are trained using a standard objective (see Equation 1);
@@
  • A standard or robust backbone classifier model combined with a front end, which is further trained using a modified variant of adversarial training (but the backbone is frozen);
@@
In that we use three types of the popular backbone models:
@@
  • ViT (Dosovitskiy et al., 2021);
A random ensemble is simply a set of three models, which are selected randomly a uniformly each time we obtain predictions $h_{\theta}(x)$ or compute gradients $\nabla_{x}L(h_{\theta}(x),y)$. It is noteworthy that randomization may unintentionally occur in PyTorch models that are inadvertently left in the training mode. In such as case, the model behaves similarly to a randomized ensemble due to enabled test-time use of dropout (Srivastava et al., 2014).
In the case of front-end enhanced models, an input image is first passed through a pre-processing network $f_{\theta_{1}}(x)$, aka the front end. Following Norouzzadeh et al. (2021), we use is a popular Gaussian Denoiser model DnCNN (Zhang et al., 2017) to process the image. DnCNN is a fully-convolutional neural network with a skip connection (see Figure 1). The front end is prepended to a standard vision backend such as ResNet (He et al., 2016) or ViT (Dosovitskiy et al., 2021), which we denote as $h_{\theta_{2}}(x)$. The full network is represented by a composite function $h_{\theta_{2}}(f_{\theta_{1}}(x))$.
Unlike many other front-end defenses (Athalye et al., 2018), DnCNN does not explicitly shatter or obfuscate gradients. Moreover, DnCNN is a fully differentiable model with a skip connection: Skip connections are commonly used to improve gradient flow (He et al., 2016).
We train the front-end enhanced model using adversarial training (see Equation 2) with a “frozen” backbone model: This means that only parameters of the front end $f_{\theta_{1}}(x)$ are being modified (using PGD to generate adversarial examples). In the case of the ResNet (and WideResNet) backends (He et al., 2016; Zagoruyko & Komodakis, 2016), we disable updates of the batch-normalization layers (Ioffe & Szegedy, 2015). Although our training procedure is somewhat similar to the approach of Norouzzadeh et al. (2021), there are two crucial differences:
@@
  • We use an extremely small (in the order of $10^{-6}$) learning rate and train for at most one epoch. In contrast, training a strong adversarial model on CIFAR datasets typically requires dozens of epochs.
As mentioned above, several models are trained using a more conventional adversarial training approach. In this case, there is no front end and there is no freezing of parameters. The model is trained for dozens of epochs using learning rates and schedules similar to those used for standard training.
@@ 4.2 Attacks
Attacks used in our work require an access to the loss function $L(h_{\theta}(x),y)$, which is used in a slightly different way for targeted and untargeted attacks. In the case of untargeted attack, the adversary needs to find a point $\hat{x}$ (in the vicinity of the clean input $x$) with the smallest “likelihood” of the correct class $y$. This entails solving the following constrained maximization problem:
@@
In the case of a targeted attack, the adversary “nudges” the model towards predicting a desired class $\hat{y}$ instead of the correct one $y$. Thus they minimize the loss of a class $\hat{y}$. This requires solving the following constrained minimization problem:
$\hat{x}=\operatorname*{arg\,min}\limits_{x\in{\mathcal{B}}_{\varepsilon}(x)}L(h_{\theta}(x),\hat{y})$ (4)
There are different approaches to solving Equation 3 or Equation 4. Most commonly, they include first-order gradient-based methods, their zero-order approximations, and random searching. In what follows we briefly survey attacks used in our work. For simplicity of exposition, we focus on the untargeted attack case.
Projected Gradient Descent (PGD)
@@
$x_{n+1}=\operatorname{proj}_{{\mathcal{B}}_{\varepsilon}(x)}\left(x_{n}+\alpha\cdot\operatorname{sign}\nabla_{{\mathbf{x}}}L(h_{\theta}(x),\hat{y})\right),$ (5)
where $\operatorname{sign}(x)$ is an element wise sign operator and $\alpha$ is a step size. PGD with restarts repeats this iterative procedure multiple time starting from different random points inside ${\mathcal{B}}_{\varepsilon}(x)$ and produces several examples. Among those, it selects a point with the maximum loss.
@@
The zero-order PGD is different from the smooth PGD in that it uses a finite-difference approximation for $\nabla_{{\mathbf{x}}}L(h_{\theta}(x),\hat{y})$. To simplify the presentation, the below description denotes $g(x)=L(h_{\theta}(x),\hat{y})$. The classic two-sided finite difference method for estimating gradients of the function $g(x)$ uses the following equation:
$\nabla_{{\mathbf{x}}}g({\bm{x}})_{i}\approx\frac{g({\bm{x}}_{0},{\bm{x}}_{1},\dots,{\bm{x}}_{i}+\alpha,\dots,{\bm{x}}_{d})-g({\bm{x}}_{0},{\bm{x}}_{1},\dots,{\bm{x}}_{i}+\alpha,\dots,{\bm{x}}_{d})}{2*\alpha}.$ (6)
@@
First we organize all variables into blocks. Then, instead of changing a single variable at a time, we add the same $\alpha$ to all variables in a block and additionally divide the estimate in Eq. (6) by the number of variables in a block. This basic procedure works well for low-dimensional CIFAR images, but not for ImageNet. For ImageNet, we first we first divide image into equal-size squares. Then, we combine squares into bigger blocks.
AutoAttack
@@
(Croce & Hein, 2020a) is a suit of attacks that has several variants of a gradient (PGD) attack as well as the query-efficient black-box SQUARE attack (Andriushchenko et al., 2020). One of the key contributions of Croce & Hein (2020b) was an Automatic Parameter-free PGD attack with two different loss functions, which are abbreviated as APGD-CE and APGD-DLR.
@@
In addition to this, APGD can be combined with an Expectation Of Transformation (EOT) algorithm (Athalye et al., 2018), which is used to approximate gradients of randomized models. EOT was shown to be effective against randomized defences (Croce & Hein, 2020a). The third algorithm is a targeted FAB attack, which was shown to be effective against gradient masking (Croce & Hein, 2020b).
@@
The standard version of AutoAttack includes all gradient attacks without EOT and the SQUARE attack. The rand version of AuttoAttack has only APGD with EOT where the the number of restarts is set to one, but the number of gradient estimation iterations is set to 20.
Attacks that bypass the front end:
@@
To circumvent the gradient-masking front end, we use two similar “removal” approaches. They both exploits the fact that a front end network is close to the identity transformation, i.e., $f_{\theta_{1}}(x)\approx x$.
@@
The first approach is a variant of the Backward Pass Differentiable Approximation (BPDA) approach with a a straight-through gradient estimator (Athalye et al., 2018). We use a previously described PGD attack and only change the way we compute gradients of the front-end enhanced network $h_{\theta_{2}}(f_{\theta_{1}}(x))$. In a general case of a straight-through gradient estimator, one would use $\nabla_{{\mathbf{z}}}h_{\theta_{2}}(z)|_{z=f_{\theta_{1}}(x)}$ instead of $\nabla_{{\mathbf{x}}}h_{\theta_{2}}(f_{\theta_{1}}(x))$. However, due to the front end $f_{\theta_{1}}(x)$ being approximately equal to the identity transformation, we use $\nabla_{{\mathbf{x}}}h_{\theta_{2}}(x)$ directly.
@@
Another approach is a variant of a transfer attack (Szegedy et al., 2014) that simply generates the adversarial example $\hat{x}$ for the network $h_{\theta_{2}}(x)$ and the clean input $x$ (again we exploit the fact that $f_{\theta_{1}}(x)\approx x$ ). To this end, we use APGD gradient attacks from the AutoAttack suit (Croce & Hein, 2020b). Then, the adversarial example $\hat{x}$ is passed to a full composite network $h_{\theta_{2}}(f_{\theta_{1}}(x))$.
@@
Table 1: Accuracy for Unprotected/Standard-training Models with a Front End (see table notes)
@@
Resnet-18 3$\times$ 92±$\pm 3$% $\pm 2$% 76±$\pm 5$% $\pm 6$%
Resnet-18 +front 3$\times$ 91±$\pm 4$% 89±$\pm 4$% 70±$\pm 6$% 74±$\pm 5$% 64±$\pm 5$% 3±$\pm 2$%
VIT-16-224 3$\times$ 97±$\pm 2$% $\pm 1$% 78±$\pm 5$% $\pm 6$%
VIT-16-224 +front 3$\times$ 96±$\pm 2$% 96±$\pm 2$% 74±$\pm 5$% 81±$\pm 4$% $\pm 6$% $\pm 1$%
@@
Resnet-18 3$\times$ 69±$\pm 5$% 5±$\pm 3$% 51±$\pm 6$% 49±$\pm 5$%
Resnet-18 +front 3$\times$ $\pm 6$% 67±$\pm 5$% 40±$\pm 6$% 54±$\pm 6$% 46±$\pm 6$% 8±$\pm 3$%
VIT-16-224 3$\times$ 85±$\pm 4$% $\pm 2$% $\pm 6$% 46±$\pm 5$%
VIT-16-224 +front 3$\times$ 86±$\pm 4$% 81±$\pm 5$% 54±$\pm 6$% 63±$\pm 5$% 45±$\pm 6$% 3±$\pm 1$%
@@
Resnet-18 3$\times$ 69±$\pm 9$% 0% $\pm 2$% 23±$\pm 8$%
Resnet-18 +front 3$\times$ 70±$\pm 8$% 68±$\pm 9$% 42±$\pm 10$% 48±$\pm 9$% 33±$\pm 9$% 0±$\pm 2$%
VIT-16-224 3$\times$ 83±$\pm 7$% 0% 18±$\pm 7$% 58±$\pm 10$%
VIT-16-224 +front 3$\times$ $\pm 8$% 80±$\pm 7$% 63±$\pm 9$% 69±$\pm 9$% 68±$\pm 9$% 0%
@@
Table 2: Summary of Results for Unprotected/Standard-training Models with a Front End
Dataset $\varepsilon$-attack Single model Best Randomized Ensembles Best RobustBench
@@
Table 3: Accuracy for Adversarially-Trained Models with a Front End on CIFAR10.
@@
5 Experiments
5.1 Experimental Setup
Evaluation Protocol
We use three datasets: CIFAR10, CIFAR100 (Krizhevsky et al., 2009), and ImageNet (Deng et al., 2009) all of which have RobustBench leaderboards for $L^{\infty}$ attacks (Croce et al., 2021). CIFAR10/100 images have a small resolution of 32 by 32 pixels, while ImageNet models use the resolution 224x224. We use attack radii that are the same as in RobustBench, namely, $\varepsilon=8/255$ for CIFAR10/CIFAR100 and $\varepsilon=4/255$ for ImageNet.
Models
For our main experiments we use four types of the models that can be classified against two dimensions:
  • Presence of the front end;
  • Single-model vs a three-model randomized ensemble.
A detailed description of these models and their training procedures is given in Section 4.1. Here we only note that for backbone types are the same between CIFAR10/CIFAR100 and ImageNet. The only exception is a WideResnet, which is used only for CIFAR10 and CIFAR100.
Attacks
We used a combination of attacks described in Section 4.2:
  • Standard 50-step PGD with 5 restarts;
  • Zero-order (finite-difference) 10-step PGD with 1 restart. For ImageNet, pixels are first grouped into 8x8 squares. For CIFAR datasets, which have very low resolution and the square size is one pixel, i.e., no such grouping is happening. Furthermore, these blocks are randomly grouped (5 squares in a group).
  • Various combinations of attacks from the AutoAttack package (Croce & Hein, 2020a). Unless specified otherwise:
    • When attacking a single (i.e., non-randomized model) we used APGD attacks without EOT (Athalye et al., 2018). Together with the SQUARE attack (Andriushchenko et al., 2020) and FAB-T (Croce & Hein, 2020b) this combination is a standard AutoAttack.
    • To attack a randomized ensemble, we use an expanded rand variant of AutoAttack. AutoAttack rand has only APGD attacks with EOT (denoted as APGD/EOT): We also use the SQUARE attack.
  • BPDA PGD with 10 steps and 5 restarts.
  • Transfer attack using AutoAttack APGD.
5.2 Discussion of Results
Summary of results
In this section we briefly review results showing that models with differentiable front ends can, indeed, appear much more robust then they actually are. In particular, when such models are assembled into a randomized ensemble, they appear to achieve near-SOTA robust accuracy virtually without degradation on clean test sets. We will then try to answer the following research questions:
  • Is this behavior due to gradient masking?
  • Is AutoAttack in principle effective against randomized ensembles?
In our experiments we focused primarily on combining standard models with a front end (except for CIFAR10). This was done to reduce experimentation cost. Indeed, training adversarially robust models for ImageNet is computationally-intensive while high-accuracy models trained using the standard objective and original/clean data are readily available.
The key outcomes of these experiments are presented in Tables 1 and 2. Table 1 includes results only for ResNet-18 (He et al., 2016) and ViT (Dosovitskiy et al., 2021). Expanded variants of Table 1, which include additional models are provided in the Appendix A.2 (Tables 6, 7, 8).
From Table 2 it can be seen that randomized ensembles, which include front-end enhanced models rival the best models on RobustBench (Croce et al., 2021), even in the case when RobustBench models were trained using additional data. At the same time, we do not observe degradation on the clean data, which indicates that a front-end should largely preserve original images. However, all front-end enhanced models have nearly zero accuracy under the BPDA PGD attack. At the same time, both the gradient attacks and the SQUARE Andriushchenko et al. (2020) attack are not particularly effective.
The SQUARE attack is quite effective against a single model without a front end. On CIFAR10 and CIFAR100 datasets (see Table 1 and Tables 6, 7 in Appendix A.2) it degrades accuracy of all models to be nearly zero. For ImageNet and ViT, the accuracy under attack is also quite low (about 20%). However, it is much less successful against randomized ensembles even when individual models are without front ends. We also find that in some cases the zero-order PGD can be more successful than other attacks: However, the attack success rate is still well below the expected 100%. To reiterate, both zero-order PGD and the SQUARE attack are particularly ineffective against randomized ensembles with front-end enhanced models.
Table 4: Accuracy under Attack for Adversarially-Trained Models on CIFAR10 for Standard vs Zero-Order PGD
@@ 5.2 Discussion of Results
What happens when an adversarially trained model is combined with a front end? We carried out a detailed investigation using a sample of CIFAR10 images and presented results in Table 3. The main difference here is that we report accuracy after each attack from the AutoAttack (Croce & Hein, 2020a) packages. Thus, we can see how successful each type of attack is.
From Table 3 we can see that adversarial training even with small $\varepsilon=2/255$, makes the model nearly “impervious” to gradient attacks when such a model is combined with a front-end. For example, a combination of all gradient attacks degrade the accuracy of a single front-end “protected” model from 95.2% to only 92.4%. At the same time the true robust accuracy of this model does not exceed 15.2%. In that, the SQUARE attack could reduce accuracy to only 41%. In the case of a randomized ensemble, though, it is not effective and the combination of attacks can reduce accuracy to only 90.8%. Curiously enough, adversarial training seems to increase resistance to both gradient and black-box attacks.
@@
It should be now clear, that the AutoAttack package (and/or its inidividual attacks) can greatly exaggerate performance of both single models and (especially) randomized ensembles. There are several indicators that this is primarily due to gradient masking. First, if we plot accuracy of the models for attacks of varying radii $\varepsilon$ (as recommended by Tramèr et al. 2020), we can see that the accuracy of adversarially-trained models without front ends quickly decreases as the attack radius $\varepsilon$ grows. In contrast, the accuracy of the front-end equipped models declines much slower and does not reach zero when $\varepsilon=1$, which is the maximum possible distortion (remember that pixel intensity ranges from 0 to 1).
@@
Second, we can see that black-box attacks, namely SQUARE and zero-order PGD, can be substantially more effective than gradient attacks. In that, for a typical adversarially trained model, the zero-order PGD is less effective (see Table 4).
@@
Third, we can see that front-end bypassing attacks, such as BPDA PGD make front-end “protected” models to be much less accurate.

### v2

@@ 3.3 Standard and Adversarial Training
To train models resistant to adversarial attacks Szegedy et al. (2014) and Goodfellow et al. (2015) proposed to expand or replace the original training set with adversarial examples generated by the model. This general idea has several concrete realizations, which are broadly termed as adversarial training. As noted by Goodfellow et al. (2015), “the adversarial training procedure can be seen as minimizing the worst-case error when the data is perturbed by an adversary.” Using a popular formalization of this statement due to Madry et al. (2018), an adversarial training objective can be written as the following minimax optimization problem:
@@
The inner maximization problem in Equation 2 corresponds to generation of adversarial examples for a given data point and a model. This problem is generally intractable: In practice, it is solved approximately using methods such as projected gradient descent (Biggio et al., 2013; Madry et al., 2018). In Section 4.2 we describe several approaches for finding adversarial examples (including PGD), which are used in our experiments.
Figure 1: Denoising front end (DnCNN) architecture proposed by Zhang et al. (2017)): Note the skip connection.
4 Models and Attacks
@@ 4.1 Model Architectures and Training
  • Standard (non-robust) models, i.e., models trained using a standard objective (see Equation 1);
@@
  • A pre-trained classifier is combined with a front end, and the resulting model is further trained using a modified variant of adversarial training, with the backbone frozen. The backbone can be either robust or standard (i.e., non-robust);
@@
In that we use three types of popular backbone models:
@@
  • ViT with varying patch sizes (Dosovitskiy et al., 2021);
A random ensemble is simply a set of three models, which are selected randomly and uniformly each time we obtain predictions $h_{\theta}(x)$ or compute gradients $\nabla_{x}L(h_{\theta}(x),y)$. It is noteworthy that randomization may unintentionally occur in PyTorch models that are inadvertently left in the training mode. In such a case, the model behaves similarly to a randomized ensemble due to enabled test-time use of dropout (Srivastava et al., 2014).
In the case of front-end enhanced models, an input image is first passed through a pre-processing network $f_{\theta_{1}}(x)$, also referred to as the front end. Following Norouzzadeh et al. (2021), we use a popular Gaussian Denoiser model DnCNN (Zhang et al., 2017) to process the image. DnCNN is a fully-convolutional neural network with a skip connection (see Figure 1). The rationale behind this architecture is to learn a residual mapping: instead of directly predicting the clean image, the network learns to estimate the noise component. This enables the model to approximate an identity transformation for clean regions while selectively correcting corrupted regions. Such a residual formulation simplifies the learning problem and accelerates convergence, especially when the corruption is relatively small.
The front end is prepended to a standard vision backend such as ResNet (He et al., 2016) or ViT (Dosovitskiy et al., 2021), which we denote as $h_{\theta_{2}}(x)$. The full network is represented by a composite function $h_{\theta_{2}}(f_{\theta_{1}}(x))$.
Unlike many other front-end defenses (Athalye et al., 2018), DnCNN does not explicitly shatter or obfuscate gradients. Moreover, DnCNN is a fully differentiable model with a skip connection, which is commonly used to improve gradient flow (He et al., 2016). Note that we use DnCNN without changing its architecture.
We train the front-end enhanced model using adversarial training (see Equation 2) with a “frozen” backbone model: This means that only parameters of the front end $f_{\theta_{1}}(x)$ are modified (using PGD to generate adversarial examples). In the case of the ResNet (and WideResNet) backends (He et al., 2016; Zagoruyko & Komodakis, 2016), we disable updates of the batch-normalization layers (Ioffe & Szegedy, 2015). Although our training procedure is somewhat similar to the approach of Norouzzadeh et al. (2021), there are two crucial differences:
@@
  • We use an extremely small (on the order of $10^{-6}$) learning rate and train for at most one epoch. Increasing the learning rate or extending training time does not yield a model that both resists strong gradient attacks and retains the clean model’s accuracy. In contrast, training a strong adversarial model on CIFAR datasets typically requires dozens of epochs.
Our training procedure also differs from that of the original DnCNN, which was trained in a fully supervised manner to remove noise. Specifically, Zhang et al. (2017) added random noise to clean images and trained the denoiser to reconstruct the original clean versions.
As mentioned above, several models in our work are trained using a more conventional adversarial training approach. In this case, there is no front end and no parameter freezing. These models are trained for dozens of epochs using learning rates and schedules similar to those used in standard training.
@@ 4.2 Attacks
Gradient attacks used in our work require access to the loss function $L(h_{\theta}(x),y)$, which is used in a slightly different way for targeted and untargeted attacks. In the case of an untargeted attack, the adversary needs to find a point $\hat{x}$ (in the vicinity of the clean input $x$) with the smallest “likelihood” of the correct class $y$. This entails solving the following constrained maximization problem:
@@
In the case of a targeted attack, the adversary “nudges” the model towards predicting a desired class $\hat{y}$ instead of the correct one $y$. Thus, they minimize the loss of a class $\hat{y}$. This requires solving the following constrained minimization problem:
[…]
There are situations in which the model’s loss surface or computational pathway causes gradient-based adversarial attacks to appear ineffective—not because the model is genuinely robust, but because the gradients are misleading, noisy, or uninformative. This phenomenon is commonly referred to as gradient masking or gradient obfuscation (Athalye et al., 2018; Tramèr et al., 2020).
There are different approaches to solving Equation 3 or Equation 4. Most commonly, they include first-order gradient-based methods, their zero-order approximations, and random search. In what follows we briefly survey attacks used in our work. For simplicity of exposition, we focus on the untargeted attack case.
Projected Gradient Descent (PGD).
@@
[…]
where $\operatorname{sign}(x)$ is an element-wise sign operator and $\alpha$ is a step size. PGD with restarts repeats this iterative procedure multiple times starting from different random points inside ${\mathcal{B}}_{\varepsilon}(x)$ and produces several examples. Among those, it selects a point with the maximum loss.
@@
[…]
$\nabla_{{\mathbf{x}}}g({\bm{x}})_{i}\approx\frac{g({\bm{x}}_{0},{\bm{x}}_{1},\dots,{\bm{x}}_{i}+\alpha,\dots,{\bm{x}}_{d})-g({\bm{x}}_{0},{\bm{x}}_{1},\dots,{\bm{x}}_{i}-\alpha,\dots,{\bm{x}}_{d})}{2*\alpha}.$ (6)
@@
First we organize all variables into blocks. Then, instead of changing a single variable at a time, we add the same $\alpha$ to all variables in a block and additionally divide the estimate in Eq. (6) by the number of variables in a block. This basic procedure works well for low-dimensional CIFAR images, but not for ImageNet. For ImageNet, we first divide an image into equal-size squares. Then, we combine squares into bigger blocks.
AutoAttack.
@@
The AutoAttack package  (Croce & Hein, 2020a)—or simply AutoAttack—is a suite of attacks that has several variants of a gradient (PGD) attack as well as the query-efficient black-box SQUARE attack (Andriushchenko et al., 2020). Note that both the SQUARE attack and the zero-order PGD attacks are query-based attacks. One of the key contributions of Croce & Hein (2020b) was an Automatic Parameter-free PGD attack with two different loss functions, which are abbreviated as APGD-CE and APGD-DLR.
@@
In addition to this, APGD can be combined with an Expectation Of Transformation (EOT) algorithm (Athalye et al., 2018), which is used to approximate gradients of randomized models. EOT was shown to be effective against randomized defenses (Croce & Hein, 2020a). The third algorithm is a targeted FAB attack, which was shown to be effective against gradient masking (Croce & Hein, 2020b).
@@
The standard version of AutoAttack includes all gradient attacks without EOT and the SQUARE attack. The rand version of AutoAttack has only APGD with EOT where the number of restarts is set to one, but the number of gradient estimation iterations is set to 20.
Transfer attack.
A transfer attack is an adversarial black-box attack where an example created to fool one (source) model is applied to a different (i.e., target) model. This exploits the phenomenon that adversarial examples often transfer across models, even when they differ in architecture, training data, or parameters (Szegedy et al., 2014).
Attacks that bypass the front end.
@@
To circumvent the gradient-masking front end, we use two similar “removal” approaches. They both exploit the fact that a front end network is close to the identity transformation, i.e., $f_{\theta_{1}}(x)\approx x$.
@@
The first approach is a variant of the Backward Pass Differentiable Approximation (BPDA) approach with a straight-through gradient estimator (Athalye et al., 2018). We use a previously described PGD attack and only change the way we compute gradients of the front-end enhanced network $h_{\theta_{2}}(f_{\theta_{1}}(x))$. In a general case of a straight-through gradient estimator, one would use $\nabla_{{\mathbf{z}}}h_{\theta_{2}}(z)|_{z=f_{\theta_{1}}(x)}$ instead of $\nabla_{{\mathbf{x}}}h_{\theta_{2}}(f_{\theta_{1}}(x))$. However, due to the front end $f_{\theta_{1}}(x)$ being approximately equal to the identity transformation, we use $\nabla_{{\mathbf{x}}}h_{\theta_{2}}(x)$ directly.
@@
Another approach is a variant of a transfer attack that simply generates the adversarial example $\hat{x}$ for the network $h_{\theta_{2}}(x)$ and input $x$. Then, the adversarial example $\hat{x}$ is passed to a full composite network $h_{\theta_{2}}(f_{\theta_{1}}(x))$. [^4] Once again, we exploit the fact that $f_{\theta_{1}}(x)\approx x$. To generate an adversarial example, we use the (gradient) APGD attack from AutoAttack (Croce & Hein, 2020b).
5 Experiments
5.1 Experimental Setup
Evaluation Protocol.
We use three datasets: CIFAR10, CIFAR100 (Krizhevsky et al., 2009), and ImageNet (Deng et al., 2009) all of which have RobustBench leaderboards for $L^{\infty}$ attacks (Croce et al., 2021). CIFAR10/100 images have a small resolution of 32 by 32 pixels, while ImageNet models use the resolution $224\times 224$. We use attack radii that are the same as in RobustBench, namely, $\varepsilon=8/255$ for CIFAR10/CIFAR100 and $\varepsilon=4/255$ for ImageNet.
Models.
For our main experiments we use four types of models that can be classified along two dimensions:
  • Presence of the front end;
  • Single-model vs a three-model randomized ensemble.
A detailed description of these models and their training procedures is given in Section 4.1. Here we only note that four backbone types are used for all three datasets. The only exception is a WideResNet, which is used only for CIFAR10 and CIFAR100.
Attacks.
We used a combination of attacks described in Section 4.2:
  • Standard 50-step PGD with 5 restarts;
  • Zero-order (finite-difference) 10-step PGD with 1 restart. For ImageNet, pixels are first grouped into 8x8 squares. For CIFAR datasets, which have very low resolution, the square size is one pixel (no such grouping occurs). Furthermore, these blocks are randomly grouped (5 squares in a group).
  • Various combinations of attacks from AutoAttack (Croce & Hein, 2020a). Unless specified otherwise:
    • When attacking a single (i.e., a non-randomized model) we used APGD attacks without EOT (Athalye et al., 2018). Together with the SQUARE attack (Andriushchenko et al., 2020) and FAB-T (Croce & Hein, 2020b) this combination is a standard AutoAttack.
    • To attack a randomized ensemble, we use an expanded rand variant of AutoAttack. AutoAttack rand has only APGD attacks with EOT (denoted as APGD/EOT). However, we also use the SQUARE attack.
  • BPDA PGD with 10 steps and 5 restarts.
  • Transfer attack using AutoAttack APGD.
@@ 5.1 Experimental Setup
Table 1: Accuracy for unprotected/standard-training models with a front end (see table notes).
@@
[…]
@@
[…]
@@
[…]
@@
Table 2: Summary of results for unprotected/standard-training models with a front end (see table notes).
Dataset $\varepsilon$-attack Single model Best Randomized Ensembles Best RobustBench (June 2025)
@@
5.2 Discussion of Results
Summary of results
In this section we briefly review results showing that models with differentiable front ends can, indeed, appear much more robust than they actually are. In particular, when such models are assembled into a randomized ensemble, they appear to achieve near-SOTA robust accuracy virtually without degradation on clean test sets. We will then try to answer the following research questions:
  • Is this behavior due to gradient masking?
  • Is AutoAttack in principle effective against randomized ensembles?
In our experiments we focused primarily on combining standard models with a front end (except for CIFAR10). This was done to reduce experimentation cost. Indeed, training adversarially robust models for ImageNet is computationally intensive while high-accuracy models trained using the standard objective and original/clean data are readily available.
The key outcomes of these experiments are presented in Tables 1 and 2. Table 1 includes results only for ResNet-18 (He et al., 2016) and ViT (Dosovitskiy et al., 2021). Expanded variants of Table 1, which include additional models are provided in the Appendix A.3 (Tables 10, 11, 12).
Using ResNet18 on CIFAR-10, we experimented with removing the skip connection from the DnCNN front-end. By design, the skip connection allows the convolutional layers to focus on learning how to suppress corruption of small magnitude, effectively acting as a near-identity transformation. Without the skip connection, training makes virtually no progress even after dozens of epochs. Interestingly, increasing the learning rate in this case leads to behavior similar to that of standard adversarial training.
Table 3: Accuracy for adversarially-trained models with a front end on CIFAR10 (see table notes).
@@ 5.2 Discussion of Results
Table 4: Accuracy under attack for adversarially-trained models on CIFAR10 for standard vs zero-Order PGD (see table notes).
@@
From Table 2 it can be seen that randomized ensembles, which include front-end enhanced models, rival the best models on RobustBench (Croce et al., 2021), even when RobustBench models were trained using additional data. At the same time, we do not observe degradation on clean data, which indicates that a front-end should largely preserve original images. However, all front-end enhanced models have nearly zero accuracy under the BPDA PGD attack. At the same time, neither the gradient attacks nor the SQUARE (Andriushchenko et al., 2020) attack are particularly effective.
@@
The SQUARE attack is quite effective against a single model without a front end. On CIFAR10 and CIFAR100 datasets (see Table 1 and Tables 10, 11 in Appendix A.3) it degrades accuracy of all models to nearly zero. For ImageNet and ViT, the accuracy under attack is also quite low (about 20%). However, it is much less successful against randomized ensembles, even when individual models are without front ends. We also find that in some cases the zero-order PGD can be more successful than other attacks. However, the attack success rate is still well below the expected 100%. To reiterate, both zero-order PGD and the SQUARE attack are particularly ineffective against randomized ensembles with front-end enhanced models.
What happens when an adversarially trained model is combined with a front end? We carried out a detailed investigation using a sample of CIFAR10 images and presented results in Table 3. The main difference here is that we report accuracy after each attack from AutoAttack (Croce & Hein, 2020a). Thus, we can see how successful each type of attack is.
From Table 3 we can see that adversarial training even with small $\varepsilon=2/255$, makes the model nearly “impervious” to gradient attacks when such a model is combined with a front-end. For example, a combination of all gradient attacks degrades the accuracy of a single front-end “protected” model from 95.2% to only 92.4%. At the same time, the true robust accuracy of this model does not exceed 15.2% (we know that the accuracy under the transfer attacks is about 15%, but it might be possible to further reduce it using some unknown approach). In contrast, the SQUARE attack could reduce accuracy to only 41%. In the case of a randomized ensemble, though, it is not effective and the combination of attacks can reduce accuracy to only 90.8%. Curiously enough, adversarial training seems to increase resistance to both gradient and black-box attacks.
@@
It should now be clear that AutoAttack (and/or its individual attacks) can greatly exaggerate performance of both single models and (especially) randomized ensembles. There are several indicators that this is primarily due to gradient masking. First, if we plot accuracy of the models for attacks of varying radius $\varepsilon$ (as recommended by Tramèr et al. 2020), we can see that the accuracy of adversarially-trained models without front ends quickly decreases as the attack radius $\varepsilon$ grows. This happens for at least two reasons:
  • Since the model is trained to resist perturbations of a radius at most $\varepsilon=8/255$, it is easily “fooled” when the attack radius increases;
  • Large perturbations can alter the underlying class semantics, so we should not expect the model to predict the initial class as the radius grows.
In contrast, the accuracy of the front-end equipped models declines noticeably slower and does not reach zero when $\varepsilon=1$, which is the maximum possible distortion (remember that pixel intensity ranges from 0 to 1).
@@
Second, we can see that black-box attacks, namely SQUARE and zero-order PGD, can be substantially more effective than standard gradient attacks. In that, for a typical adversarially trained model, the zero-order PGD is less effective (see Table 4). This suggests that the gradients provide little useful information about the true optimization direction—a characteristic phenomenon of gradient masking.
@@
Third, we can see that front-end bypassing attacks—such as BPDA PGD—make front-end “protected” models much less accurate. The ineffectiveness of standard gradient-based attacks, contrasted with the success of stronger approaches such as transfer attacks, is a well-recognized indicator of gradient masking Tramèr et al. (2020).